# Gravitational waves from strong first order phase transitions

### v1

@@ Acknowledgments
JC acknowledges useful and fruitful discussions with members of the Computational Field Theory group of Helsinki, including Lauri Niemi, Lorenzo Giombi, Deanna Hooper, Jani Dahl. We thank Jani Dahl for comments on the manuscript. All data collected for this manuscript with the exception of material in Appendix E was produced using resources from CSC IT Center for Science via project 2008297 (codenamed GRAPE). Earlier stages of this project (with smaller grid sizes) used projects 2003695 and hy5787 on the same machine. We acknowledge PRACE for previously awarding access to the HAZEL-HEN facility in Germany at the High-Performance Computing Center Stuttgart (HLRS), where the material from Appendix E was used. JC (ORCID ID $0$-0002-3375-0997) acknowledges support from Research Council Finland grant 354572 and ERC grant CoCoS 101142449; MH (ORCID ID $0$-0002-9307-437X) from Academy of Finland grant 333609, Research Council of Finland grant 363676 and STFC grant ST/X000796/1; KR (ORCID ID $0$-0003-2266-4716) from the ERC grant CoCoS 101142449 and the Research Council of Finland grant 354572; and DJW (ORCID ID $0$-0001-6986-0517) from Research Council of Finland grants 324882, 328958, 349865 and 353131.
@@ Appendix A Advection routines
One of the update operations, $\dot{W}=-\partial_{i}(Wv^{i})$, where $W$ is a state variable, such as $E,Z_{j}$, supports high, low, or hybrid version of flux computation. State variables are defined as cell-centered or edge-centred. Velocities are face-centered. The flux of a cell-centred variable through face $i$ is $Wv^{i}$. The rate of change of a cell-centred state variable at site $k$ is
@@
The low-order version of flux computation corresponds to donor cell advection [80]. For donor cell, if the flux is positive, $W_{[{k+1/2}]}=W_{[k]}$, or if it is negative, $W_{[{k+1/2}]}=W_{[k+1]}$.
For the higher-order scheme, a function $\Phi(r_{k})$ is introduced, where $r_{k}=(W_{[}k]-W_{[k-1]})/(W_{[k+1]}-W_{[k]})$ is a ratio of successive gradients direction $i$, and the flux in direction $i$ defined as
$\begin{split}&f(W_{[k+1/2]})\\ &=f^{L}(W_{[k+1/2]})\\ &-\Phi(r_{k})\bigg{(}f^{L}(W_{[k+1/2]})-f^{H}(W_{[k-1/2]})\bigg{)}\end{split}$ (84)
@@
$\Phi(r)=\left\{\begin{array}[]{cl}\displaystyle\max[0,\min(1,r)],&\text{(MinMod)},\\ \displaystyle\frac{r+|r|}{1+|r|},&\text{(Van Leer)}.\end{array}\right.$ (85)
In our case, we have observed that for the velocity flows set up by the phase transitions we simulated, spurious oscillations were still present in the compressional mode with either Van Leer or MinMod limiters. These were not only visible when rendering visualization slices, but also on velocity power spectra, shown in 7. This can be seen in Fig. 7 for a smaller detonation test run with volume $2048^{3}$ where a “bump” appears in both higher-order compressional velocity power spectra at roughly $kR_{*}\approx 2\times 10^{2}$, which is not present with the low-order scheme.
Figure 7: The compressional and vortical velocity power spectra $\mathcal{P}_{v_{\parallel},\perp}$ (in shades of green and pink, respectively) for three smaller volume ($2048^{3}$) test runs of the detonation case ($v_{\text{w}}=0.92$, $\alpha_{\text{n}}=0.67$), with three different choice of flux limiter. From top to bottom: donor cell, MidMod, and Van Leer (see discussion around Eq. 85). The runs are initialised with $N_{b}=64$ bubbles, giving mean bubble spacing $R_{*}=512$. The bubbles have the same positions, and the times plotted are $t=2800,3200,3600,4000,4400T_{c}^{-1}$ in all three. We additionally overlay a shock power spectra fit (shown in dashed black, see Eq. 87) to a high-$kR_{*}$ region of the donor cell plot.
@@
To further justify our choice, and demonstrate that the downturn in the velocity power spectrum away from a power law at high wavenumber in donor cell advection has a physical explanation, we fit the analytical formula for a power spectrum dominated by shocks from [34]. The following behaviour is expected at higher wavenumbers,
@@
where $\mathcal{A}$, $n$, and $\kappa_{s}$ are free parameters, $z=kR_{*}$ and $\mathcal{M}$ is a modulating function given by,
@@
The fitting results for $kR_{*}\in[101.3,1609.3]$ are presented in Fig. 7 (top) for the latest velocity power spectrum. The fit itself is excellent and we report the following best-fit parameters: $\mathcal{A}=0.26$, $n=-1.43$ and $\kappa_{s}=3.33\times 10^{2}$. The parameter $\kappa_{s}$ gives the shock width, $\delta_{s}=R_{*}/\kappa_{s}\simeq 1.53/T_{\text{c}}$.
@@ Appendix B Energy evolution
Figure 8: The behaviour of volume averaged energies in each test case, with the top panel corresponding to the detonation ($v_{\text{w}}=0.92$, $\alpha_{\text{n}}=0.67$) and the lower one to the deflagration ($v_{\text{w}}=0.44$, $\alpha_{\text{n}}=0.5$). We present the total fluid-scalar energy in $e_{tot}$, the thermal energy of the fluid the $e_{Q}$, the fluid kinetic energy $e_{K}$, the trace energy $e_{\theta}$, the scalar field potential and gradient energies in $e_{V}$ and $e_{D}$.
@@ Appendix C Fluid profile from one bubble simulations
We present here the results of 1 bubble 3d simulations, where a single bubble is placed in a smaller lattice ($1024^{3}$) and allowed to collide with itself around the periodic volume. All other parameters are the same as the full $4096^{3}$ grid simulations, as is the effective $R*$ (bubble centre spacing). The resulting fluid profiles at position $(x,y=0,z=0)$ before and after the fluid shells interact are shown in Fig. 9, along with the ideal hydrodynamic solutions for the expanding bubbles. These simulations indicate how close the fluid profiles are to their asymptotic form when they collide.
@@
Figure 9: The 3-velocity and weighted 4-velocity fluid profiles measured from 3d one bubble simulations at $1024^{3}$ lattice size before the collision of the fluid shells. The times displayed correspond to $t=500T_{c}^{-1}$ and for the detonation (left-hand-side panel), and $t=740T_{c}^{-1}$ for the deflagration (right-hand-side panel).
@@ Appendix D Derivation of GW kernels
In the integrations, we assume that $\omega_{AB}\bar{t}\gg 1$, which is the case as wavenumbers greater than about $10/R_{*}$ and times $\bar{t}\gtrsim R_{*}$. Then the limits on $s$ can be taken to infinity, giving
@@
Figure 10: Comparison of numerical and theoretical gravitational wave spectra for weak phase transitions. We show the instantaneous rate of change of the gravitational wave power spectra $\dot{\mathcal{P}}_{\text{gw}}$ computed via the convolution of velocity power spectra of compressional modes, $C_{\Delta}[{v_{\parallel}},{v_{\parallel}}]$, and estimated from simulation data with a numerical difference. Both are given in dimensionless form, normalised by a factor of $H_{\text{n}}^{2}R_{*}$. Left: detonation ($\alpha_{\text{n}}=0.0046$, $v_{\text{w}}=0.92$), with differences taken between times $1480/T_{c}$ and $2080/T_{c}$. Right: deflagration ($\alpha_{\text{n}}=0.0046$, $v_{\text{w}}=0.44$) using times $3920/T_{c}$ and $5600/T_{c}$.
@@ Appendix E Weak phase transitions and sound shell model prediction
In this appendix we compare previous data for weak phase transitions ($\alpha_{\text{n}}=0.0046$) [30] to the predictions from the velocity convolution formula 68 and to the sound shell model [36, 37], as realised in the PTtools package (see Fig. 10). Wall speeds are $v_{\text{w}}=0.92$ and $v_{\text{w}}=0.44$. More precisely, we compare the growth rate of the spectrum, as computed from a numerical difference $\Delta\mathcal{P}_{\text{gw}}(k)\Delta t$ between spectra at times given in the caption, scaled by $H_{\text{n}}^{2}R_{*}$. The sound shell model computation is carried out at the default precision, and with the simultaneous nucleation option. Excellent agreement is found.

### v2

@@ Acknowledgments
JC acknowledges useful and fruitful discussions with members of the Computational Field Theory group of Helsinki, including Lauri Niemi, Lorenzo Giombi, Deanna Hooper, Jani Dahl. We thank Jani Dahl for comments on the manuscript. All data collected for this manuscript with the exception of material in Appendix E was produced using resources from CSC IT Center for Science via project 2008297 (codenamed GRAPE). Earlier stages of this project (with smaller grid sizes) used projects 2003695 and hy5787 on the same machine. We acknowledge PRACE for previously awarding access to the HAZEL-HEN facility in Germany at the High-Performance Computing Center Stuttgart (HLRS), where the material from Appendix E was used. JC (ORCID ID 0000-0002-3375-0997) acknowledges support from Research Council Finland grant 354572 and ERC grant CoCoS 101142449; MH (ORCID ID 0000-0002-9307-437X) from Academy of Finland grant 333609, Research Council of Finland grant 363676 and STFC grant ST/X000796/1; KR (ORCID ID 0000-0003-2266-4716) from the ERC grant CoCoS 101142449 and the Research Council of Finland grant 354572; and DJW (ORCID ID 0000-0001-6986-0517) from Research Council of Finland grants 324882, 328958, 349865 and 353131.
@@ Appendix A Advection routines
Our implementation of one of the update operations, $\dot{W}=-\partial_{i}(Wv^{i})$, where $W$ is a state variable, such as $E$ or $Z_{j}$, supports high-order, low-order, and hybrid versions of the flux computation. State variables are defined as cell-centered or edge-centred. Velocities such as the 3-velocity $v^{i}$ are generally face-centered. The flux $f(W)$ of a cell-centred variable through face $i$ is $Wv^{i}$. The rate of change of a cell-centred state variable at site $k$ is
@@
The low-order version of flux computation (denoted $f^{L}$) corresponds to donor cell advection [32]. For donor cell, if the flux is positive, $W_{[{k+1/2}]}=W_{[k]}$, or if it is negative, $W_{[{k+1/2}]}=W_{[k+1]}$.
For the higher-order scheme (where the flux is now denoted $f^{H}$), a function $\Phi(r_{k})$ is introduced, where $r_{k}=(W_{[k]}-W_{[k-1]})/(W_{[k+1]}-W_{[k]})$ is a ratio of successive gradients in direction $i$, and the flux in direction $i$ defined as
$f(W_{[k+1/2]})\\ =f^{L}(W_{[k+1/2]})-\Phi(r_{k})\bigg(f^{L}(W_{[k+1/2]})-f^{H}(W_{[k+1/2]})\bigg).$ (83)
@@
$\Phi(r)=\begin{cases}\max[0,\min(1,r)],&\text{(MinMod)},\\[10.0pt] {\displaystyle\frac{r+|r|}{1+|r|}},&\text{(Van Leer)}.\end{cases}$ (84)
Figure 7: The compressional and vortical velocity power spectra $\mathcal{P}_{v_{\parallel},\perp}$ (in shades of green and pink, respectively) for three smaller volume ($2048^{3}$) test runs of the detonation case ($v_{\text{w}}=0.92$, $\alpha_{\text{n}}=0.67$), with three different choice of flux limiter. From top to bottom: donor cell, MidMod, and Van Leer (see discussion around Eq. 84). The runs are initialised with $N_{b}=64$ bubbles, giving mean bubble spacing $R_{*}=512$. The bubbles have the same positions, and the times plotted are $t=2800,3200,3600,4000,4400T_{c}^{-1}$ in all three. We additionally overlay a shock power spectra fit (shown in dashed black, see Eq. 86) to the high-$kR_{*}$ region of the donor cell plot.
In our case, we have observed that for the velocity flows set up by the phase transitions we simulated, spurious oscillations were still present in the compressional mode with either Van Leer or MinMod limiters. These were not only visible when rendering visualization slices, but also on velocity power spectra, shown in Fig. 7.
This figures show the results of test runs with the detonation simulation parameters (albeit at smaller volume $2048^{3}$). At roughly $kR_{*}\approx 2\times 10^{2}$, a ‘bump’ appears in the compressional velocity power spectra obtained with the higher-order schemes, which is not present with the low-order scheme.
@@
To further justify our choice, and demonstrate that the downturn in the velocity power spectrum away from a power law at high wavenumber in donor cell advection has a physical explanation, we fit the analytical formula for a power spectrum dominated by shocks from Ref. [40]. The following behaviour is expected at higher wavenumbers,
@@
where $\mathcal{A}$, $n$, and $\kappa_{s}$ are free parameters, $z=kR_{*}$ and $\mathcal{M}$ is a modulating function given by
@@
The fitting results for $kR_{*}\in[101.3,1609.3]$ are presented in Fig. 7 (top) for the latest velocity power spectrum. The fit itself is excellent and we report the following best-fit parameters: $\mathcal{A}=0.26$, n=1.43n=-1.43 and $\kappa_{s}=3.33\times 10^{2}$. The parameter $\kappa_{s}$ gives the shock width, $\delta_{s}=R_{*}/\kappa_{s}\simeq 1.53/T_{\text{c}}$.
@@ Appendix B Energy evolution
Figure 8: The behaviour of volume averaged energies in each test case, with the left panel corresponding to the detonation ($v_{\text{w}}=0.92$, $\alpha_{\text{n}}=0.67$) and the right one to the deflagration ($v_{\text{w}}=0.44$, $\alpha_{\text{n}}=0.5$). We present the total fluid-scalar energy in $e_{tot}$, the thermal energy of the fluid the $e_{Q}$, the fluid kinetic energy $e_{K}$, the trace energy $e_{\theta}$, the scalar field potential and gradient energies in $e_{V}$ and $e_{D}$.
@@ Appendix C Fluid profile from one bubble simulations
We present here the results of one-bubble 3d simulations, where a single bubble is placed in a smaller lattice ($1024^{3}$) and allowed to collide with itself around the periodic volume. All other parameters are the same as the full $4096^{3}$ grid simulations, as is the effective RR* (bubble centre spacing). The resulting fluid profiles at position $(x,y=0,z=0)$ before the fluid shells interact are shown in Fig. 9, along with the ideal hydrodynamic solutions for the expanding bubbles. These simulations indicate how close the fluid profiles are to their asymptotic form when they collide. We remark that the intial acceleration phase when the fluid relaxes towards its self-similar solution implies that computing $r/t$ using total simulation time will mildly underpredict $r/t$. As a result we use linear fitting of $r$ to $t$ to calculate the time offset $t_{0}$ and present the corrected measure in Fig. 9. There we can see that the detonation fluid shell is very close to its asymptotic profile, while the deflagration fluid shell is still undergoing acceleration. To further support this observation we additionally plot the wall velocity estimates ($v_{w,b}$ and $v_{w,\phi}$ defined in Section III) measured in the 3-dimensional one bubble simulations.
@@
Figure 9: The 3-velocity and weighted 4-velocity fluid profiles measured from 3d one bubble simulations at $1024^{3}$ lattice size before the collision of the fluid shells. The times displayed correspond to $t=500T_{c}^{-1}$ and for the detonation (left-hand-side panel), and $t=740T_{c}^{-1}$ for the deflagration (right-hand-side panel). Note that the time variable ($r/t$) is offset by time $t_{0}$ spent in the initial acceleration phase (ie. $r/(t-t_{0})$). For the detonation $t_{0}=8.46T_{c}^{-1}$ and for the deflagration $t_{0}=32.50T_{c}^{-1}$. We also plot the wall velocity estimates $v_{w,b}$ and $v_{w,\phi}$.
@@ Appendix D Derivation of GW kernels
In the integrations, we assume that $\omega_{AB}\bar{t}\gg 1$, which is the case as wavenumbers greater than about $10/R_{*}$ and times $\bar{t}\gtrsim R_{*}$. Then, the limits on $s$ can be taken to infinity, giving
@@
Figure 10: Comparison of numerical and theoretical gravitational wave spectra for weak phase transitions. We show the instantaneous rate of change of the gravitational wave power spectra $\dot{\mathcal{P}}_{\text{gw}}$ computed via the convolution of velocity power spectra of compressional modes, $C_{\Delta}[{v_{\parallel}},{v_{\parallel}}]$, and estimated from simulation data with a numerical difference. Both are given in dimensionless form, normalised by a factor of $H_{\text{n}}^{2}R_{*}$. Left: detonation ($\alpha_{\text{n}}=0.0046$, $v_{\text{w}}=0.92$), with differences taken between times $2400/T_{c}$ and $2800/T_{c}$. Right: deflagration ($\alpha_{\text{n}}=0.0046$, $v_{\text{w}}=0.44$) using times $3800/T_{c}$ and $4200/T_{c}$.
@@ Appendix E Weak phase transitions and sound shell model prediction
In this appendix we compare previous data for weak phase transitions ($\alpha_{\text{n}}=0.0046$) [53] to the predictions from the velocity convolution formula 69 and to the sound shell model [55, 49], as realised in the PTtools package (see Fig. 10). Wall speeds are $v_{\text{w}}=0.92$ and $v_{\text{w}}=0.44$. More precisely, we compare the growth rate of the spectrum, as computed from a numerical difference $\Delta\mathcal{P}_{\text{gw}}(k)\Delta t$ between spectra at times given in the caption, scaled by $H_{\text{n}}^{2}R_{*}$. The sound shell model computation is carried out at the default precision, and with the simultaneous nucleation option. Excellent agreement is found.